OpenStudy (anonymous):

The function C(r) = 2pie-r gives the circumference of a circle of a radius 'r'. Find the circumference of a circle with radius 3.5inches. Give the exact value as a multiple of pie and an approximate value, rounded to the nearest hundredth.

OpenStudy (aravindg):

this ones too easy we have C(r)=2*pi*r here r=3 C(3)=2*pi*3 =6*pi

OpenStudy (aravindg):

now pi=3.14 approximately

hero (hero):

@ChristaB You should write your function as \[C(r) = 2\pi r\]

OpenStudy (aravindg):

substituting c(3)=6*3.14 =18.84

OpenStudy (aravindg):

did that help?

OpenStudy (anonymous):

isn't the 'r' = to 3.5?

OpenStudy (aravindg):

oops ...typo ..yes its 3.5

OpenStudy (aravindg):

sorry

OpenStudy (aravindg):

the answer will be 7*pi=21.98

OpenStudy (anonymous):

so it will look like this: 3.5 (\[7 \] + ???

OpenStudy (aravindg):

it will be 2*3.5*pi

OpenStudy (aravindg):

2*3.5=7

OpenStudy (aravindg):

gt it ?

OpenStudy (anonymous):

so, punching the numbers into the calculator, using the pi button.. the answer I came up with is 21.9911..... is this correct?

OpenStudy (aravindg):

i gt 21.98

OpenStudy (anonymous):

Now, to round that to the nearest hundredth.... the answer becomes 22.00???

OpenStudy (aravindg):

yepe

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

I think so... my confusion right now though is the rounding to the nearest hundredth... does the 21.99114 rounding to the nearest hundredth become 22.